You have assigned the following values to these three firms: Price Upcoming Dividend Growth Beta US Bancorp $ 29.80 $ 3.20 8.80 % 1.59 Praxair 59.15 1.51 13.00 2.20 Eastman Kodak 36.50 1.00 11.50 0.98 Assume that the market portfolio will earn 15.50 percent and the risk-free rate is 5.70 percent. Compute the required return for each company using both CAPM and the constant-growth model. (Do not round intermediate calculations and round your final answers to 2 decimal places.)
Answer #1

Solution:

The required return as per CAPM is kp = rf + beta (rm - rf)

Where:

kp is the required return , rf is the risk free rate, rm -> market return , beta -> beta value of the firm

As per Constant growth model we have:

V = D1 / (k - g)

Where, V is the value of the stock, g -> annual growth rate of the dividends , k -> required rate of return, D1 ->next year's dividend.

Given rf = 5.70% , rm = 15.50%

Company Price Upcoming Div Growth Beta CAPM Constant Growth
US Bancorp 29.80 3.20 8.80% 1.59

kp = 5.70% + 1.59(15.50% - 5.70%)

= 5.70% + 1.59(9.8%)

= 5.70% + 15.582%

kp =21.282%

v = D1 / k -g

29.80 = 3.20/ k - 8.80%

k - .088 = 3.20/29.80

k - 0.088 = .10738255

k = .19538255

= 19.53%

Praxair 59.15 1.51 13% 2.20

kp = 5.70% + 2.20(15.50% - 5.70%)

= 5.70% + 2.20(9.8%)

= 5.70% + 21.56%

kp =27.26%

59.15 = 1.51/ k - 13%

k - 13% = 1.51/59.15

k - 0.13 = .025528318

k = .155528318

= 15.55%

Eastman Kodak 36.50 1 11.50% .98

kp = 5.70% + .98(15.50% - 5.70%)

= 5.70% + .98(9.8%)

= 5.70% + 9.604%

kp =15.304%

36.50 = 1/ k - 11.50%

k - 11.50% = 1/36.50

k - 0.115 = .02739726

k = .14239726

= 14.23%